### v1

# Insecurity Through Obscurity: Veiled Vulnerabilities in Closed-Source Contracts

Abstract.
Most blockchains cannot hide the binary code of programs (i.e., smart contracts) running on them. To conceal proprietary business logic and to potentially deter attacks, many smart contracts are closed-source and employ layers of obfuscation. However, we demonstrate that such obfuscation can obscure critical vulnerabilities rather than enhance security, a phenomenon we term insecurity through obscurity. To systematically analyze these risks on a large scale, we present skanf, a novel EVM bytecode analysis tool tailored for closed-source and obfuscated contracts. skanf combines control-flow deobfuscation, symbolic execution, and concolic execution based on historical transactions to identify and exploit asset management vulnerabilities. Our evaluation on real-world Maximal Extractable Value (MEV) bots reveals that skanf detects vulnerabilities in 1,028 contracts and successfully generates exploits for 373 of them, with potential losses exceeding \$9.0M. Additionally, we uncover 40 real-world MEV bot attacks that collectively resulted in \$900K in losses.
1. Introduction
Smart contracts (Wood et al., 2014) are autonomous programs on a blockchain, offering strong integrity and availability guarantees. However, because most blockchains lack privacy, smart contract bytecode is publicly visible. While transparency is a key feature, it becomes a drawback when a smart contract’s logic embodies competitive advantages or proprietary strategies.
To mitigate this, a variety of obfuscation techniques are commonly applied. A prominent example is the Maximal Extractable Value (MEV) bots, which are optimized smart contracts deployed by MEV searchers to perform arbitrary, sandwich attacks and other on-chain actions to extract MEV (Daian et al., 2020). Six of the top 10 MEV bots, ranked by the number of transaction bundles they sent (libMEV, 2025), employ some form of obfuscation.
There are several plausible reasons to use obfuscation, including to keep proprietary strategies private and to deter attacks (i.e., security from obscurity). However, our analysis of the recent “Destroyer Inu” incident (detailed in Sec. 3) reveals that obfuscation can also introduce insecurity. In this case (detailed in Sec. 3), 22 ETH (worth \$51,056 at the time) were stolen from a closed-source [^1] A smart contract is closed-source if its source (e.g., Solidity code) is not publicly available. Binary code is always publicly available on Ethereum. smart contract equipped with layers of obfuscation. The attack exploited a well-known vulnerability—tx.origin phishing (by Example, 2025)—that standard analysis tools can readily detect. But because the contract was obfuscated, the vulnerability went unnoticed until it was exploited. In this case, obscurity arguably renders the contract less secure by hindering analysis, making this a prime instance of insecurity through obscurity (Meunier, 2008) [^2] Compare this with modern cryptographic schemes following Kerckhoffs’s principle that a system’s security should not be compromised if adversaries uncover its method of operation (Petitcolas, 2019), or, in the words Shannon: “the enemy knows the system” (Shannon, 1949)..
This work.   Due to the lack of effective analysis tools, little is known about the security of closed-source smart contracts, especially those that handle large volumes of assets and present lucrative targets for attackers. In this paper, we aim to develop effective code analysis techniques to bridge this gap.
We focus on a broad class of security vulnerabilities that we refer to as asset management vulnerabilities, issues that allow an adversary to manipulate a smart contract’s logic to steal assets (i.e., transferring tokens from the victim to the attacker herself). While our techniques are general, we use MEV bots as a real-world testbed to evaluate the efficacy of our tool and as concrete examples to study the security of closed-source smart contracts in the wild.
Our goals can be summarized as the following three research questions (RQs):
  • RQ1: How to effectively deobfuscate the control flow given smart contract bytecode? Vulnerability detection is only possible if this can be accomplished.
  • RQ2: How to effectively detect asset management vulnerabilities in closed-source smart contracts? Even if obfuscation is removed, detecting such vulnerabilities can be challenging, given the complexity of MEV bots. Existing tools often timeout when applied to MEV bots, even without obfuscation.
  • RQ3: How many MEV bot contracts have been exploited in practice, and how much did they lose? Answering this question helps in understanding the severity of asset management vulnerabilities and estimating how much loss could be reduced if skanf were used by searchers.
We answer these questions using the following methods. First, we analyze notable attacks on closed-source contracts, and distill the vulnerabilities that allowed them and the methods used to execute them. Then, we use our new-found understanding to devise a novel deobfuscator that successfully uncovers crucial parts of closed-source contracts which are opaque to established tools. Making use of our deobfuscator and insights, we build a vulnerability inspector and an automatic attack generator. We combine our deobfuscator, vulnerability inspector and attack generator into a tool we call skanf  [^3] The name skanf stands for Sen, Kaihua, Aviv ’n Fan.. By applying skanf to closed-source contracts we collect, we count the number of closed-source contracts that have potential security vulnerabilities and automatically synthesize the exploits against them. We further estimate a lower bound of the amount of funds that can be taken from them using historical Ethereum data.
1.1. Challenges
Numerous tools (Dedaub, 2025; ConsenSys, 2025; Frank et al., 2020; Gritti et al., 2023; Grech et al., 2019) are available to identify and exploit smart contract bugs. To illustrate the challenges in answering our research questions, we highlight three core limitations of existing tools that render them inefficient for our tasks.
Control-flow obfuscation.   Mainstream languages like Solidity and Vyper rely on jump tables with static jump destinations to manage function dispatch and control flow. However, contracts may employ “indirect jumps” to make the control flow dependent on runtime values to achieve obfuscation.
To the best of our knowledge, no existing tools effectively handle control flow obfuscation. State-of-the-art static analysis tools, such as Gigahorse (Grech et al., 2019, 2022; Lagouvardos et al., 2024), classify basic blocks as unreachable when a contract relies on indirect jumps to determine control flow. Decompilation tools such as Dedaub, which claims to “successfully decompile over 99.98% of deployed contracts on the Ethereum blockchain” (Dedaub, 2025), fail to function effectively when facing control flow obfuscation. For instance, Dedaub can only analyze 4% of code in the top-2 obfuscated MEV bots. Symbolic execution tools like Mythril (ConsenSys, 2025), ETHBMC (Frank et al., 2020), and Greed (Gritti et al., 2023; Ruaro et al., 2024) raise errors when jump destinations are symbolic.
Lack of fine-grained analysis.   Deciding asset management vulnerabilities requires a fine-grained analysis of a smart contract’s logic, and simple pattern matching as done by existing tools does not suffice. For example, JACKAL (Gritti et al., 2023) flags CALL instructions with a fixed function selector as safe, which overlooks vulnerabilities where the attacker manipulates function parameters (instead of the function selector). As another example, one might flag tx.origin-based authentication as insecure (because it appears similar to tx.origin phishing (by Example, 2025)), but later steps may rectify the problem.
Performance due to complex logic.   In addition to asset management, the contracts we focus on also involve trading, flash loans, swaps on different DEXs, and other operations and interactions. This complexity leads to performance issues in static analysis, as the tools must analyze many control-flow paths and resolve intricate data-flow dependencies, increasing both processing time and memory usage. Moreover, the rapid growth of execution paths and the complexity of constraints make symbolic execution inefficient due to path explosion and the high computational cost of constraint solving. While some tools support concolic execution (ConsenSys, 2025), generating high-quality seed input remains a manual process that does not scale.
1.2. Our Methods
Recovering control flow.   We propose a method to de-obfuscate control flows exploiting the following EVM feature: Unlike traditional instruction sets such as x86-64 or ARM64, where jump instructions can target any executable address, the EVM requires all valid jump destinations to be explicitly marked with a JUMPDEST instruction (Wood et al., 2014). This allows us to determine the boundaries of basic blocks directly from the bytecode and reconstruct jump tables, even when jump destinations are determined dynamically at runtime. Our tool can instrument the bytecode with a reconstructed jump table to significantly improve the code coverage of subsequent analysis, which is crucial for efficacy.
Detecting asset management vulnerabilities.   To not rely on a specific pattern for detecting asset management vulnerabilities, we broadly look for vulnerable calls in the bytecode of a smart contract that an adversary can trigger with adversarial parameters. As a naive example, if a smart contract exposes a CALL instruction where the adversary can set the target function as well as its input, the adversary can call WETH.transfer with the “to” parameter set to itself to drain the victim’s WETH. Many attacks, however, do not require the adversary to control the entirety of the inputs.
Our approach draws inspiration from traditional program vulnerability detection, where the combination of symbolic execution and taint analysis has been used to address similar problems (Schwartz et al., 2010). To detect asset management vulnerability, we first use symbolic execution to identify execution paths leading to a CALL statement. Then, we apply taint analysis to examine whether the inputs to the CALL statement can be controlled by an adversary. Attacks are possible even if the adversary cannot control the entirety of the inputs, so we carefully enumerate cases where an adversary can manipulate critical parameters, such as the recipient address for approve or transfer.
Concolic execution based on historical transactions.   As mentioned above, a naive application of symbolic execution does not scale to handle real-world smart contracts of interest. Concolic execution solves the performance issues of symbolic execution by combining concrete and symbolic execution to mitigate path explosion (Sen, 2007). By leveraging concrete seed inputs to guide path exploration and simplify constraint solving, concolic execution can reduce the search space, improving efficiency. The “quality” of the seed input is critical, but there are no general algorithms to generate high-quality ones.
We observe, however, that high-quality inputs are already readily available on the blockchain: all historical transactions are publicly visible, and many of them invoke CALL instructions through various paths. This allows us to automatically extract high-quality seed inputs from external and internal transactions recorded on the blockchain, significantly improving runtime performance and the efficacy of vulnerability detection.
1.3. Implementation and Evaluation of skanf
Putting these ideas together, we developed skanf based on Gigahorse (Grech et al., 2019) and Greed (Gritti et al., 2023; Ruaro et al., 2024). At a high level, skanf takes the bytecode and the historical transactions of a given smart contract as the input and outputs exploits if the smart contract is vulnerable.
The workflow of skanf consists of three stages. First, it performs static analysis to detect indirect jumps and attempts to deobfuscate the control flow by inserting a switch table that converts all indirect jumps into direct jumps. Second, skanf parses the contract’s historical transactions as concrete seed input for concolic execution and identifies asset management vulnerabilities during the execution. For each identified vulnerability, skanf synthesizes exploits using constraint-solving, with adversary-controlled parameters — such as the ERC-20 token address and transfer amount — as constraints. The exploit is then validated and reported.
Evaluation results.   To evaluate the efficacy of skanf against real-world smart contracts, we compile a dataset of 6,554 MEV bot contracts and assess skanf alongside three state-of-the-art tools. For each tool, we limit the execution time to 10 minutes and count the number of vulnerable smart contracts it can detect. Our evaluation shows that skanf outperforms the other tools: it detects 1,028 vulnerable smart contracts and successfully generates exploits for 373 of them. If exploited, the potential losses of these vulnerabilities would exceed 9.0 million USD. In comparison, other tools such as ETHBMC (Frank et al., 2020) and JACKAL (Gritti et al., 2023) detected 0 and 18 vulnerable contracts, respectively. While Mythril identified 574 vulnerable contracts, our verification confirmed that 84% of them are false positives.
To understand if skanf can flag attacks that already happened, we conduct an empirical study on MEV bot attacks in the wild. We discovered 40 exploits targeting 12 MEV bot contracts, resulting in a total loss of \$900K. This confirms that asset management vulnerabilities pose a serious threat to bot security. All exploited contracts are flagged by skanf—had skanf been available to developers, these exploits could have been detected early on. Even after these attacks occurred, only three had been previously reported, probably because no tool was available to automatically scan for such exploits.
Contributions
In short, we make the following contributions:
  • We propose a novel deobfuscation method to remove control flow obfuscation and enable further analysis.
  • We propose a concolic execution approach that uses historical transactions of a contract as concrete inputs to identify potential vulnerabilities in smart contracts.
  • Building on these two techniques, we implement skanf. We plan to release the code publicly.
  • We evaluate the efficacy of skanf against real-world closed-source and obfuscated smart contracts. skanf identifies 1,028 vulnerable contracts and automatically exploits 373 of them. The potential losses of these vulnerabilities exceed \$9.0M.
  • We discover 40 real-world attacks targeting MEV bots, which we categorize as MEV phishing attacks. These attacks have resulted in an estimated loss of \$900K to searchers. All of the attacked smart contracts are flagged by skanf.
Responsible Disclosure.   As our exploits can readily compromise existing smart contracts, we responsibly disclosed them to the affected parties through a dedicated blockchain messaging service (Blockscan Chat by Etherscan (Etherscan Team, 2023)), following the same disclosure practice adopted by prior research (Qin et al., 2023). Moreover, we adhere to several additional responsible practices. We test all exploits on a local replication of the Ethereum blockchain that is disconnected from the public network. Unless the attacks have already occurred in the wild and the vulnerabilities are publicly known, we do not disclose the specific addresses of vulnerable contracts in the paper.
2. Preliminaries
2.1. Ethereum and Smart Contracts
Ethereum.   Ethereum is the second-most popular blockchain, and the foremost smart contract-enabled blockchain. It relies on a Proof-of-Stake (PoS) mechanism, wherein actors can become part of the set of validators operating the blockchain by staking (i.e., depositing) Ethereum’s native token, ETH. Ethereum’s mechanism sequentially assign one validator to propose a block, i.e., a batch of transactions, every 12 seconds. The validator chosen to act as a proposer has a free hand in constructing its block.
MEV supply chain.   Today’s Ethereum block construction pipeline involves several specialized entities. At one end, we have searchers who identify profitable opportunities, e.g., transactions that can be front-run (Zhou et al., 2023), and assemble transaction bundles that exploit them. These bundles are then sent to builders, who specialize in constructing profitable blocks, and who compete in an auction-esque process to have their blocks chosen by upcoming proposers.
Smart contract.   Smart contracts are computer programs stored and executed on the blockchain. Once deployed, contracts can be invoked and interacted with using transactions. Ethereum contracts are specified in EVM bytecode, but can be written using high-level languages like Solidity (Team, 2025b) and then compiled for deployment (Wood et al., 2014). Besides using high-level languages, sophisticated developers can further optimize contracts–such as reducing gas costs–by directly writing bytecode or using low-level languages like Huff (Team, 2025a). While Ethereum smart contracts are stored on-chain using low-level EVM bytecode, developers can choose to open-source their contracts’ high-level code on platforms like Etherscan (Etherscan, 2025a).
The Application Binary Interface (ABI) specifies how to encode function calls and decode responses when interacting with a smart contract on Ethereum. If a contract is open-sourced, its ABI becomes publicly accessible; otherwise, the ABI is unavailable, making it difficult for the public to know how to interact with the contract.
The EVM is the execution environment for Ethereum smart contracts. Specifically, smart contracts execute in a stack-based virtual machine, meaning operations are performed by manipulating data on a last-in-first-out (LIFO) stack rather than via registers. Contracts are specified using low-level EVM bytecode composed of opcodes such as CALL, JUMP, and JUMPI. Control flow within the EVM is managed through the JUMP and JUMPI instructions, which provide unconditional and conditional branching, respectively.
Indirect jump.   In the EVM, an indirect jump occurs when a jump instruction (JUMP/JUMPI) is executed and the top stack value (i.e., the jump destination) is not a statically known constant, but rather a runtime-computed value. Obfuscation techniques often rely on indirect jumps to hide control flow, making it difficult to statically determine which path may be executed.
External call.   During execution, a smart contract can invoke another contract or externally owned account (EOA) using one of four opcodes: CALL, CALLCODE, DELEGATECALL, and STATICCALL. These opcodes enable the caller to pass along Ether (value), execution gas, and input data (calldata), facilitating inter-contract interactions. They differ in how the execution context is propagated, particularly with respect to msg.sender, tx.origin, and storage access. In this paper, we focus on the CALL instruction, which underlies common patterns such as calling an ERC-20 token’s transfer function; we refer to such invocations as external calls.
The EVM provides two global variables to track the source of a call: tx.origin and msg.sender (accessed via ORIGIN and CALLER opcodes). The variable tx.origin always refers to the original EOA that initiated the transaction, remaining constant across the entire call chain. In contrast, msg.sender refers to the immediate caller of the current function, which may be an EOA or another contract, and it changes with each external call. For clarity, we refer to tx.origin as origin and msg.sender as caller throughout the rest of this paper.
We refer to the target contract that an external call aims to invoke as the target address. The calldata consists of two parts: the first four bytes, named function selector, which identify the target function to be executed in the target contract; and the remaining bytes, which specify the arguments passed to that function.
Asset management.   In this paper, we focus on ERC-20 token assets held by smart contracts. The ERC-20 standard defines two methods for transferring tokens (Ethereum Foundation, 2025). An account can invoke transfer on an ERC-20 token contract to send tokens directly to another account. Alternatively, it can use approve to authorize another account to transfer tokens from its balance using transferFrom.
2.2. Threat Model
Our adversary model captures the ability of a typical attacker targeting smart contracts. They can create transactions, deploy smart contracts, and interact with deployed contracts but cannot access the private keys of other accounts, breach the integrity of the blockchain, or manipulate the block-building process.
Asset management vulnerability.   Our focus is on asset management vulnerabilities, i.e., vulnerabilities that would allow an adversary to transfer tokens from the contract’s account to the adversary’s. In particular, weak adversaries such as our own can do so when a smart contract relies on obfuscation to hide its lack of access control checks. Another possibility is if the smart contract verifies that a token transfer is allowed by checking if the invoking transaction originated from a white-listed address; then, adversaries can employ phishing attacks wherein searchers attempt to interact with a seemingly innocent contract, which then performs the malicious transfer on the adversaries’ behalf.
3. Case Study: the Destroyer Inu Attack
A notable attack on a closed-source smart contract that relies on external inputs for its operation took place on July 1st, ’24, causing a loss of 22 ETH (worth \$51,056 at the time) (0xprincess, 2024). As the attack involves an attacker-created ERC-20 token called “Destroyer Inu”, we use the same name for the attack itself. This case study shows that control-flow obfuscation adopted by the victim contract ends up hiding known vulnerabilities from analysis tools—without stopping real-world attackers—arguably making contracts less secure.
tx.origin phishing attack.   The core issue is a classic vulnerability where tx.origin is compared to a hard-coded trusted address, but hidden behind control flow obfuscation. This vulnerability is known as the tx.origin phishing attack (by Example, 2025), dating back to at least 2016 (PeterBorah, 2016). In such attacks, an attacker tricks the victim into calling a malicious contract, which then calls the victim’s contract; since both calls have the same tx.origin, the malicious smart contract can execute the victim smart contract with the same privilege as the victim herself.
1assembly {
2entry 0x0:
3  let from := origin() // tx.origin
4 let expected := 0xdead...beef
5  if iszero(eq(from, expected)) {revert(0, 0)} // access check
6 // read two bytes from calldata as the jump destination.
7 let data := calldataload(0x84) // suppose it’s 0x0a00...
8 let dst := shr(0xf0, data)
9 jump dst // jump 0x0a00
10
11jumpdest 0x0a00:
12  let token := calldataload(0x86)
13  let to := calldataload(0xa6)
14  let value := calldataload(0xc6)
15  let ptr := mload(0x40)
16  mstore(ptr, 0xa9059cbb) // ERC-20 transfer
17 mstore(add(ptr, 4), to)
18 mstore(add(ptr, 36), value)
19 // call target.transfer(to, value)
20 let success := call(gas(), token, 0, ptr, 68, 0, 0)
21 if iszero(success) { revert(0,0) }
22
23jumpdest 0x0b00: ...
24}
Figure 1. Pseudo-Yul MEV bot smart contract. The code starting at 0x0a00 contains a vulnerable function call whose input is determined by the calldata, allowing an adversary to transfer any amount of ERC-20 tokens from the MEV bot’s account to their own. Note that jump and jumpdest are not part of Yul syntax but are included for better readability.
[…]
tx.origin phishing attack, obfuscated.   In the Destroyer Inu Attack, the victim contract is obfuscated. We use pseudo-Yul code to illustrate the smart contract’s logic in Fig. 1. First, at the entry point of the MEV bot’s smart contract in lines 3 to 5, it implements an access control mechanism based on the transaction’s origin to prevent unauthorized addresses from interacting with the contract. Then, in lines 7 to 9, the contract proceeds by jumping to a specific place in code, as determined by the calldata supplied by the transaction. This is unlike “standard” smart contracts that match function calls using the so-called function selector defined as the first 4 bytes of the calldata (Contributors, 2025). In this MEV bot smart contract specifically, when the two calldata bytes at positions 0x85 and 0x86 are 0x0a00, the program jumps to the code segment starting at jumpdest 0x0a00. The functionality of this segment is to manage the transfer of a specific amount of an ERC-20 token to a target address. Therefore, it loads the address of the ERC-20 token, the recipient address, and the amount from the calldata (lines 12 to 14). Then, in lines 15 to 21, it constructs the input based on these parameters and invokes the ERC-20 transfer.
Exploit.   To an attacker who can remove control-flow obfuscation, the vulnerability is immediate: the victim contract allows transferring any amount of any ERC-20 tokens held by this contract to another address if an attacker can bypass the tx.origin check via many forms of tx.origin phishing, which is precisely what happened. We provide additional details on the phishing attack in Appx. A.
This case study suggests that obfuscation can hide vulnerabilities and motivates us to develop skanf to systematically evaluate asset management vulnerabilities in closed-source smart contracts.
4. The design of skanf
4.1. Overview
Figure 2. Overview of skanf. The input consists of the smart contract’s bytecode and historical transactions, while the output is the verified exploit. In the first stage (ⓐ), we identify the obfuscated smart contracts (➀) and recover their control flow information (➁). In the second stage (ⓑ), skanf parses historical transactions as inputs for concolic execution (➂). Then it identifies potential vulnerabilities within the contract and labels adversary-controllable parameters (➃). Finally, in the last stage (ⓒ), skanf generates potential exploits for the identified vulnerabilities (➄), which are then validated in ➅.
[…]
The inputs to skanf are the bytecode and historical transactions of a given smart contract; if an exploit is found, skanf outputs a verified exploit against the contract. Fig. 2 provides an overview of our system, which consists of three stages:
  • Control flow deobfuscation. The first stage identifies and removes control flow obfuscation. The output of this stage is bytecode with a reconstructed switch table to remove indirect jumps that prevent code analysis tools from functioning properly.
  • Concolic execution powered by historical transactions. Smart contracts of interest have complex logic and dependencies, requiring concolic execution to speed up vulnerability discovery. Unlike tools that rely on manual input, skanf can automatically identify high-quality concrete data from historical transactions that enable efficient exploration. The output of this stage is a list of potential vulnerabilities, i.e., insufficiently protected invocations of CALL that could be exploited to steal assets from the smart contract.
  • Exploit generation and validation. To exploit a vulnerability identified in the previous stage, one must craft a transaction that triggers the vulnerable CALL with specific adversarial inputs. A challenge is to ensure the execution of code after the said vulnerable CALL to succeed without errors.
4.2. Control Flow Deobfuscation
In the first stage, we aim to deobfuscate the target smart contract to facilitate further analysis. Technically, we need to identify indirect jumps (i.e., whose destinations are determined at runtime) and convert them to direct jumps (i.e., whose destinations are statically specified) to make the control flow explicit and analyzable using standard static techniques.
Identifying indirect jumps.   We first perform static analysis on the bytecode to identify indirect jumps, a common sign of control flow obfuscation. In practice, the destination of an indirect jump is typically based on dynamic inputs such as calldata. To identify such jumps, we extract all jump instructions (JUMP and JUMPI) and build the control flow graph. We then analyze whether the destination of each jump depends on dynamic inputs (i.e., calldata, value, memory, and storage). Each indirect jump is recorded in the set $\mathcal{J}=\{(j_{1},d_{1}),\dots,(j_{n},d_{n})\}$ where $j_{i}$ is the PC of the jump instruction and $d_{i}$ is the stack variable that determines the jump destination at runtime (e.g., v58, v158 in Fig. 3).
If $\mathcal{J}$ is empty, the contract does not use control flow obfuscation, and skanf proceeds to the second stage. Otherwise, we proceed to deobfuscate the indirect jumps.
Figure 3. Assembly code for the obfuscated smart contract and how to rewrite the indirect jump as a direct jump by inserting a crafted jump table.
[…]
Deobfuscation.   The EVM spec requires that every valid jump destination must be explicitly marked with a JUMPDEST instruction. This property allows us to statically scan the disassembled bytecode, collect all legal jump destinations, and instrument the bytecode with a switch table that replaces indirect jumps with direct jumps to restore analyzable control flow.
The idea is best illustrated with an example. Suppose the original smart contract contains an indirect jump to an address stored in variable v. Most static analysis tools struggle to handle this since there are too many possible branches. Our idea is to rewrite the code as follows:
1if v == 0x0a00 : jump 0x0a00
2else if v == 0x0b00 : jump 0x0b00
3else if v == ...
where 0x0a00, 0x0b00, and so on are all jump destinations extracted from the bytecode. The number of branches is exactly the number of valid jump destinations. This transformation preserves the original control flow while ensuring that all jump destinations are immediate constants encoded in the bytecode.
Replacing every jump instruction with the above code snippet is inefficient, so we reuse the above logic by adding a switch table. A switch table is a structure commonly used to implement multi-way branching based on a runtime value. We implement the switch table as a sequence of conditional checks. Each entry compares a runtime value — previously used as a jump destination — with a statically known case value. If a match is found, the program performs a direct jump to that case’s target. Otherwise, it proceeds to the next entry. We insert the switch tables at PC 0xe000 because the smart contract has a size limit of 24 KB (0x6000 bytes) (Buterin, 2016) — ensuring that the added code does not interfere with existing logic.
If the original jump instruction is JUMPI, special handling is required because it involves two stack values: the jump destination (on top of the stack) and the condition (just below it). To preserve the original semantics when redirecting execution to the switch table, we must maintain the correct stack layout. Before inserting the new jump destination, we insert a SWAP1 to move the condition to the top of the stack. This ensures that when the rewritten JUMPI executes, it evaluates the condition correctly while using the switch table entry as the destination. We also insert a POP immediately after JUMPI to remove the unused destination value from the stack in case the condition is false and the jump does not occur, thereby avoiding stack imbalance.
A complete example is shown in Fig. 3. For every indirect jump instruction at PC $j\in\mathcal{J}$, we rewrite the bytecode to push 0xe000 onto the stack. This redirects execution to the start of the switch table. Each switch table entry is implemented as a short code snippet. For instance, consider the entry at PC 0xe0010xe009. It begins by duplicating the runtime value for comparison. Then, it pushes a known destination (e.g., 0x0a00) and performs an equality check using the EQ instruction. If the check passes, the program jumps to an intermediate gadget at PC 0xf0000xf005, which cleans up the stack and performs the final jump to 0x0a00. This intermediate step is necessary because the comparison only consumes the duplicated value, leaving the original runtime value on the stack. If the comparison fails, execution continues to the next table entry (e.g., at PC 0xe00a), which repeats the process for another possible destination. This continues until a matching entry is found.
Note that the switch table includes all possible jump destinations within the contract. Although some of these destinations may never be used, our evaluation Sec. 5.2 indicates that this approach does not introduce significant overhead.
After this transformation, all jump targets in the contract are encoded as immediate values, and the control flow becomes fully explicit and statically analyzable.
4.3. Vulnerability Detection
In this stage, we identify potential asset management vulnerabilities in the smart contract — specifically, opportunities for an adversary to steal assets. We define a vulnerability oracle that flags any vulnerable CALL instruction as one that is (1) reachable from a public entry point and (2) accepts inputs that are either adversary-controlled or fixed but risky (e.g., a known ERC-20 token address).
Note that a vulnerability does not guarantee a successful exploit. Akin to traditional vulnerabilities like buffer overflows, where a flaw alone does not guarantee arbitrary code execution, as additional conditions are often required. In this paper, we define an exploit as a transaction that transfers assets from the victim contract to the adversary’s account.
We detect vulnerabilities in this stage and generate exploits in the next. To this end, we first extract concrete seed inputs from historical transactions in order to perform concolic execution, thereby speeding up the exploration process. Additionally, we perform taint analysis during the path exploration leading to a vulnerable CALL instruction. The taint analysis results are then used to determine whether the reachable CALL instruction is one of the vulnerabilities of interest.
Extracting seed input from historical transactions.   A key challenge in analyzing real-world smart contracts lies in their complex logic, especially when involving multiple CALL instructions to interact with other smart contracts, such as DEXs, NFTs, and lending protocols. To efficiently explore such code, we adopt concolic execution (Sen, 2007), a hybrid analysis technique that uses concrete seed inputs to guide path exploration and simplify constraint solving, thereby reducing the search space and improving efficiency.
In our setting, an ideal seed input is a set of parameters that trigger an external call to transfer assets out of the smart contract. This prioritizes the exploration of “high-risk” paths involving asset transfers. Crafting seed inputs for closed-source smart contracts is hard because the ABI (Application Binary Interface) is unknown. Randomly generated seed inputs often fail to reach vulnerable paths. Fortunately, high-quality seed inputs are already readily available on-chain in historical transactions. Given a target contract, we retrieve the associated transactions that are executed successfully and extract the relevant inputs, which serve as concrete seeds for concolic execution.
Focusing on asset-transferring paths may introduce false negatives, since not all potentially vulnerable paths are explored. However, this trade-off is acceptable — discovering even a single asset management vulnerability suffices to demonstrate that the contract’s assets are at risk. As shown in Sec. 5.3, this approach allows us to uncover more vulnerabilities than existing tools.
Specifically, we collect historical transactions of a given contract $C$ and extract external calls that involve ERC-20 transfers by inspecting standardized transfer event logs (Ethereum Foundation, 2025). For each relevant transaction, we extract all associated external calls, keeping only those where the target is $C$. We further filter out calls with empty calldata (e.g., plain ETH transfers), as they typically do not involve asset management logic.
Concolic execution.   From the parsed historical transactions, we extract all external calls to the target contract, including the origin, caller, calldata, and value. These, along with the corresponding block number (to reflect the correct blockchain state), are used as concrete inputs for concolic execution.
While concolic execution can help identify reachable CALL instructions, we still need to analyze whether the inputs to CALL instructions are vulnerable. During concolic execution, we perform on-the-fly taint propagation to track how tainted calldata bytes, i.e., taint sources representing adversarial inputs, propagate to critical input to CALL instructions. Specifically, we break down the input to the CALL instruction into different parameters, including the target address, the target function (function selector), and individual function arguments. Taint propagation begins when calldata is loaded (e.g., via CALLDATALOAD or CALLDATACOPY) and continues dynamically through the EVM’s stack, memory, and storage. We conservatively propagate taint through arithmetic, logical, and copy operations. The taint sinks we monitor are external calls, where we check whether the target address, target function, or any function arguments depend on tainted data.
When concolic execution encounters a CALL, it performs two key operations. First, we inspect the instruction’s parameters to determine whether it represents a potential asset management vulnerability. Then, based on taint analysis, we selectively replace calldata bytes that influence the CALL parameters with symbolic variables, while preserving concrete values for all other bytes. In stage ⓒ, we generate exploits by restarting concolic execution with these modified calldata bytes, which allows us to construct an exploit based on the values specified by the adversary.
Identifying asset management vulnerability.   After the taint analysis, skanf identifies CALL parameters that can be determined by the adversary via their input. We refer to such parameters as adversary-controllable. However, even when some parameters are fixed (thus cannot be changed by an adversary), the CALL instruction may still be vulnerable. For example, in Fig. 1, the function selector (0xa9059cbb) is fixed, yet the adversary still controls the target address of the call and can construct the calldata to transfer all WETH to their account. Therefore, we also analyze fixed parameters to assess whether they pose a risk.
In the context of asset management vulnerabilities, a fixed target address may still be problematic if it corresponds to an ERC-20 token contract (e.g., WETH). Similarly, a fixed function selector may be risky if it maps to functions like transfer or approve. We refer to such parameters as risky, since they can still lead to asset loss even without full adversarial control.
Once we determine that both the target address and function selector of the vulnerable CALL instruction are either adversary-controllable or risky, we examine whether bytes 5 to 36 of its calldata — which are interpreted as the first argument if the function is transfer or approve — can be arbitrarily set by the adversary, noting that if the function selector is controllable, the adversary can choose transfer or approve. If this is the case, we classify the corresponding CALL instruction as a potential asset management vulnerability. Additionally, we record whether the second argument (i.e., the transfer amount) is fixed or under adversarial control.
Fallback mode.   If concolic execution does not identify any vulnerabilities, we fall back to symbolic execution, which performs a broader search for all potentially vulnerable CALL instructions and determines whether they are reachable. Specifically, we first enumerate all CALL instructions in the contract. For each, we use the CFG and call graph to check whether there exists a path from a public function to the CALL. If such a path is found, we initiate symbolic execution from the corresponding public function using a symbolic byte string as calldata.
As symbolic execution progresses, we keep track of symbolic expressions for all values in the EVM stack, memory, and storage. These symbolic values are updated with each instruction to reflect how they depend on input variables. When the execution encounters a conditional branch (i.e., JUMPI), we fork the execution into two paths — one for each possible outcome — and add the corresponding condition to the path constraints. To manage complexity, we prune any infeasible path. This can happen in two cases: either the accumulated constraints become unsatisfiable (i.e., there’s no input that could follow that path), or the CFG shows that the path cannot reach the target CALL instruction. Handling access control checks — like those involving msg.sender or tx.origin — can be particularly tricky. To improve our chances of passing these checks, we try two configurations. First, we set both the caller and the origin to a predefined adversary address. If that fails (for instance, if the contract requires a specific tx.origin), we try again with the origin set to the one recorded from a previously observed transaction during concolic execution.
Once symbolic execution discovers a feasible path to a CALL instruction, we analyze it using the same criteria as in concolic execution — checking whether any of its parameters are adversary-controllable or otherwise risky — to determine whether it constitutes a vulnerability.
Filtering out false positives.   Symbolic execution may introduce false positives due to its imprecise modeling of external calls and incomplete representations of storage or execution context (Gritti et al., 2023). For instance, a CALL instruction may appear reachable in symbolic analysis but cannot be triggered under any concrete execution. To eliminate such cases, we perform additional validation for each identified vulnerability.
This validation is done using a local blockchain instance to simulate a concrete execution environment. We supply the concrete calldata–solved from any symbolic constraints–as well as the caller and origin addresses to verify whether the CALL instruction is actually executed. Unlike exploit generation, we do not require the transaction to succeed or result in asset transfer; we only check whether the vulnerable CALL is triggered. This is sufficient to confirm the vulnerability’s existence. Exploitability, which requires additional constraints to be satisfied, is addressed in the next stage.
At the end of this stage, skanf reports a confirmed vulnerability (an example output is illustrated in Fig. 4), recording the PC of the CALL instruction, the risky or adversary-controllable parameters, and other contextual information such as the caller, origin, block height, and the concrete calldata used. These details are forwarded to the next stage for exploit generation.
1{
2 "caller": "0xdead...beef",
3 "origin": "0xdead...beef",
4 "blockNumber": 20000000,
5 "callPC": "0xac5", // PC of the vulnerable CALL
6 "calldata": "12345678SSSS...", // SSSS represents the symbolic data
7 "targetAddress": "*", // target address is adversary-controllable
8 "functionSelector": "0xa9059cbb", // fixed but risky
9 "destination": "*",
10 "amount": "*"
11}
Figure 4. Example output of skanf for a vulnerability.
[…]
4.4. Exploit Generation and Validation
At the final stage, skanf attempts to synthesize and validate a concrete exploit for each identified vulnerability.
Automatic exploit generation.   Given a vulnerability identified in the previous stage, skanf attempts to construct a transaction that successfully exploits it. The transaction is built at the specific block height where the vulnerability is observed. The “to” address is set to the victim contract. The “gas” and “gas price” fields are configured to exceed the minimum requirements to ensure execution and inclusion in a block. For the “value” field, if any value is acceptable, we set it to zero; otherwise, we use the required amount. The “from” address is set to either the adversary’s address (if no constraints are specified) or to a specific origin address if one was recorded during vulnerability detection. The latter case is typically caused by tx.origin-based authorization, which can be bypassed through phishing attacks, as is the case with the Destroyer Inu attack. Given how easy such phishing attacks are in practice (see  Sec. 6 for real-world evidence), we conservatively assume that the attacker can always set the “from” value to the required one.
The key step in exploit generation is constructing the correct calldata to ensure the transaction executes successfully. Recall from stage ⓑ that for each vulnerability, we record which CALL parameters — the target address, function selector, and arguments — are adversary-controllable. If the target address is controllable, we constrain it to the address of an ERC-20 token held by the victim contract with a non-zero balance. By repeating the attack with different addresses, we can systematically target each token. Similarly, if the function selector and arguments are controllable, we constrain them to values that mimic a realistic attack: we set the selector to transfer, the recipient (to) to the adversary’s address. For the transfer amount, we query the token balance of the victim contract at the specified block height and set this balance as the transfer amount, aiming to steal the maximum possible amount as an adversary would.
According to the EVM specification (Wood et al., 2014), a transaction is considered successful only if execution halts at a valid stop instruction such as STOP or RETURN. To ensure this condition is met, we resume concolic execution from the vulnerable CALL and symbolically execute forward until a stop instruction is reached. This introduces additional path constraints, which are solved to adjust the synthesized calldata and complete exploit generation.
Exploit validation.   Finally, skanf validates the exploit in a local execution environment to ensure that it executes successfully and produces the expected behavior.
To do this, we simulate the blockchain environment at the block height specified in the exploit. This includes loading the contract code, account balances, and persistent storage state. We then execute the synthesized transaction in this environment.
After execution, we check whether the transaction succeeded. We further assess exploit success by inspecting the emitted event logs: for ERC-20 exploits, we verify the presence of a Transfer or Approval event consistent with the expected asset movement, following the ERC-20 standard (Ethereum Foundation, 2025). If both checks pass, skanf confirms the exploit as valid.
5. Evaluation
5.1. Experimental Setup
Implementation.   We implement skanf on top of Gigahorse (Grech et al., 2019, 2022) and Greed (Gritti et al., 2023; Ruaro et al., 2024), comprising 2.2K lines of Python code and 210 lines of Soufflé code. In more technical detail, we develop a plugin for Gigahorse to analyze bytecode and recover the control flow of obfuscated contracts. We then use Gigahorse to lift the (deobfuscated) bytecode to its register-based Intermediate Representation (IR). The vulnerability detection and exploit generation are implemented based on Greed, which attempts to automatically synthesize exploits through constraint solving given the following inputs: the register-based IR of a contract along with the CFG and call graph generated by Gigahorse. For a given contract, we fetch its recent transactions using the Etherscan API (Etherscan, 2025a). To validate exploits, we simulate their execution using the Python implementation of the EVM (Foundation, 2025; ethpwn, 2025).
Dataset.   We use real-world MEV bots as the target to evaluate the efficacy and performance of skanf. We first compile a dataset of the addresses of a total of 6,554 MEV bots from multiple sources, including existing works (Yang et al., 2024a; Öz et al., 2024; Heimbach et al., 2024), public dashboards (libMEV, 2025; Etherscan, 2025b), and results from MEV analysis tools such as MEV inspect (Flashbots, 2024).
Testbed.   We perform our experiments on Ubuntu 22.04 LTS, with an Intel Xeon Platinum 8380 (80 cores), 128GB RAM and 8TB SSD.
5.2. RQ1: Deobfuscation Effectiveness
The effectiveness of skanf in addressing control-flow obfuscation is measured using the code coverage metric. Specifically, a tool’s coverage of a given smart contract is the fraction of reachable code blocks. Low code coverage indicates that a significant portion of the smart contract cannot be analyzed by the tool.
We compare Gigahorse’s coverage with and without deobfuscation for all smart contracts in our dataset. To zoom in on real-world obfuscation practices used by major actors, we repeat the comparison on the top 10 and top 100 most active MEV bots since the Merge (ranked by the number of bundles they sent), as reported by the “libMEV” online dashboard (libMEV, 2025). We choose these bots because sophisticated MEV searchers who remain active are strongly motivated to keep their business logic private, making them more likely to adopt control flow obfuscation in their code.
Figure 5. Cumulative Distribution Function (CDF) of the code coverage of smart contracts in our dataset for both Gigahorse and our tool, skanf. As can be observed, skanf’s CDF is concentrated near 100%, while that of Gigahorse is spread out and is consistently higher (i.e., worse) than skanf, indicating that skanf outperforms Gigahorse.
[…]
Fig. 5 presents the cumulative distribution function (CDF) of the code coverage for the top 10, top 100, and all smart contracts in the dataset, respectively. Dashed lines represent Gigahorse, and solid lines represent skanf. We observe that among the top 10 MEV bots, six are highly obfuscated (with code coverage below 10%), and skanf successfully increases the code coverage to 100% for five of them. Note that skanf fails on one bot contract because it relies on the preliminary output from Gigahorse for bytecode analysis. Although Gigahorse generally produces preliminary output even for highly obfuscated contracts, in this particular case, it fails to do so, which prevents skanf from proceeding. Comparing the code coverage of three MEV bot groups (top 10, top 100, and all bots), we find that control flow obfuscation is more prevalent among the most active MEV bots, underscoring the importance of our approach.
For all bots, we observe that for 5% of the smart contracts, the original code coverage is below 50%. In contrast, fewer than 0.4% of smart contracts have deobfuscated code coverage below 50%. This comparison highlights the effectiveness of skanf in mitigating obfuscation. For the small fraction (0.4%) of contracts where deobfuscation does not significantly improve code coverage, our manual analysis reveals two reasons: either Gigahorse fails to produce any preliminary output before deobfuscation, or some blocks within the contracts appear to be unreachable, rather than hidden by control flow obfuscation.
Answer to RQ-1 skanf effectively addresses control flow obfuscation in smart contracts, for 90% of contracts with an initial coverage below 50%, we can successfully increase their code coverage.
5.3. RQ2: Vulnerability Detection Effectiveness
We now evaluate the effectiveness of skanf in detecting vulnerabilities in smart contracts, as measured by the number of identified vulnerabilities. To benchmark our tool, we conduct a comparative evaluation against three state-of-the-art bytecode-based symbolic execution tools: Mythril (commit: 9e9ee39(ConsenSys, 2025), ETHBMC (commit: e887f33(Frank et al., 2020), and JACKAL (commit: 3993e5c(Gritti et al., 2023). In the comparative evaluation, we run each tool for a maximum of 10 minutes per contract and impose a memory limit of 20 GB.
5.3.1. Vulnerability detection by skanf
To illustrate how concolic execution enhances the effectiveness of vulnerability detection, we conduct an ablation study by running skanf in full symbolic execution mode without any concrete inputs, which we refer to as the baseline mode. In our evaluation, skanf detects 719 vulnerable smart contracts in the baseline mode. When executed in concolic mode, the number of detected vulnerable smart contracts increases to 1,028. This highlights a significant improvement in vulnerability detection achieved through concolic execution, which simplifies constraint solving and prioritizes exploration of the vulnerable CALL instruction within the limited time.
Comparison with state-of-the-art tools.   Mythril alerts when it detects a publicly reachable CALL instruction where the adversary can control the target address of the CALL. In the evaluation, Mythril flags 574 potentially vulnerable smart contracts. To filter out false positives–similar to our approach for skanf–we validate the vulnerabilities identified by Mythril by executing the exploit transaction it generates in a local blockchain instance. Specifically, we check whether the vulnerable CALL instruction is triggered by the adversary and whether the adversary can control other parameters to steal the contract’s assets.
In the end, we find that only 89 of the flagged contracts are true positives (skanf identifies 1,028). Among these, 68 are also covered by skanf. For the remaining contracts not identified by skanf, we manually inspect them to understand the cause. We find that these cases are not covered because identifying vulnerabilities requires an accurate CFG, and we build the contract’s CFG using Gigahorse, which fails to generate CFGs effectively for these contracts (e.g., due to out-of-memory errors during bytecode lifting).
ETHBMC also raises an alert when it detects a CALL where the adversary controls the target address and transfer amount (Frank et al., 2020; Gritti et al., 2023). Note that this captures only a subset of asset management vulnerabilities, as an adversary can still steal assets even if the target and amount are fixed, as long as they control the recipient.
However, in our evaluation, ETHBMC did not detect any vulnerabilities. Our further analysis shows that low code coverage and timeouts are the two main reasons. For about 36% of the contracts, ETHBMC achieved less than 25% code coverage during execution. Additionally, for approximately 27% of the contracts, ETHBMC encountered timeout errors while solving constraints. These findings indicate limitations in ETHBMC’s ability to analyze obfuscated smart contracts and identify asset management vulnerabilities in contracts with complex logic.
JACKAL detects reachable CALL instructions where the adversary can control the target address and the function selector of the CALL. For the same reason as with ETHBMC, this pattern represents a subset of asset management vulnerabilities. In our evaluation, JACKAL detects 18 vulnerable contracts, all of which are included in the set of 1,028 vulnerable contracts we identified.
5.3.2. Runtime of skanf
Beyond its effectiveness in identifying vulnerabilities, we further compare skanf against the other three state-of-the-art symbolic execution tools in terms of runtime.
Figure 6. CDF of the time taken by skanf, Mythril, ETHBMC, and JACKAL to analyze the smart contract.
[…]
As shown in Fig. 6, we observe that skanf achieves shorter analysis time than the other tools we evaluated. In concolic mode, it can analyze about 90% of the contracts in the dataset within 100 seconds, whereas other tools, such as Mythril, complete the analysis for less than 60% within the same time limit. Although both skanf and JACKAL are built on Gigahorse and Greed, skanf optimizes analysis by halting exploration once a vulnerability is found, which may cause false negatives. In contrast, JACKAL continues exploring. We consider this tradeoff acceptable, as one asset management vulnerability already signals risk to the contract’s assets.
5.3.3. Exploit generation and loss estimation
For each smart contract with identified vulnerabilities, we generate and validate exploits. We compute the loss if these exploits were realized.
In total, we successfully exploited 373 of 1,028 vulnerable contracts. We validated all 373 exploits locally. For the remaining contracts, we manually inspected 10 samples and confirmed that they were not false positives (i.e., the vulnerability is valid but the synthesized exploits failed). In some cases, the synthesized exploits fail because the contracts contain additional CALL instructions requiring further parameter refinement. We leave this for future work, and skanf already significantly outperforms existing tools.
We further calculate a lower bound on the potential loss if the synthesized exploits were carried out. Since the balance of the victim contract changes over time, we use two reference points: the current balances as of the time of writing (March 8th, 2025), or the historical high balances (including vulnerable historical versions; see Appx. B), which represent a worst-case attack. For simplicity of comparison, we use the token price on March 8th, 2025, to convert the loss to USD. We focus on seven major ERC-20 tokens: WETH, WBTC, USDC, USDT, DAI, UNI, and LINK, so our results are a lower bound. If a contract’s vulnerability is limited to specific ERC-20 tokens, we only consider the affected tokens.
The results are shown in Tab. 1. The “current” and “maximal” columns refer to the current balances and historical high balances, respectively. As shown in the table, the total estimated economic loss caused by the synthesized exploits amounts to approximately 9.0 million USD. Among them, WETH accounts for the largest portion, with an estimated loss of approximately 6.4 million USD. The significant loss suggests that these profitable MEV bots are vulnerable and exposed to substantial financial risks despite their high earnings.
Table 1. A conservative estimate of the potential loss if the synthesized exploits were carried out.
Token Current Maximal
Amount Loss (\$) Amount Loss (\$)
WETH 139.3 306,639.0 2,895.1 6,374,070.6
WBTC 0.2 17,003.1 20.0 1,723,308.9
USDC 29,982.5 29,982.5 296,716.3 296,716.3
USDT 23,789.5 23,789.5 282,295.5 282,295.5
DAI 41,802.7 41,802.7 185,609.3 185,609.3
UNI 13,647.3 97,620.7 14,699.8 105,149.0
LINK 1,573.0 23,988.0 3,061.2 46,682.6
Total 540,825.5 9,013,832.2
Answer to RQ-2 skanf identifies 1,028 potentially vulnerable MEV bot contracts and successfully generates effective exploits for 373 of them, with a potential loss exceeding \$9.0M. Its vulnerability detection significantly outperforms other tools.
6. Attacks in the Wild
Evaluation in Sec. 5 confirms that skanf can identify new vulnerabilities not reported or exploited in practice. In this section, we present an empirical study of realized attacks exploiting asset management vulnerabilities in the wild. The goal of this study is threefold. First, we aim to identify realized attacks that have not been detected or reported to gain an understanding of the prevalence of real-world exploits against closed-source smart contracts. Second, we run skanf on the exploited smart contracts and evaluate whether skanf can discover vulnerabilities in them, from which we quantify the potential loss that could have been saved by skanf. Finally, as discussed in Sec. 4.3, we claim that attackers can control the originator of the transaction to bypass the origin check. We seek to validate this assumption against real-world incidents.
6.1. Detecting Attacks
We first need to collect real-world attacks to support the analysis objectives stated earlier. However, we emphasize that our goal is not to conduct a comprehensive measurement study (in particular, we do not aim to be exhaustive). Therefore, we use a relatively simple method that targets low false positives.
We introduce two heuristics to detect exploits of asset management vulnerabilities in MEV bot contracts. First, based on how MEV bots work, we observe that a contract never called by an MEV bot is unlikely to call the MEV bot—in other words, external calls from a “strange” smart contract that an MEV bot had no prior interaction with are highly likely to be attack attempts (also illustrated in Fig. 8, Appx. A). The second rule is that the malicious smart contract should not be an ERC-20 or ERC-721 token that strictly follows the ERC-20/ERC-721 standards; this filters out false positives. Finally, we manually verify whether the assets were transferred to the adversary accounts. The adversary accounts can be distinguished from the searchers’ own accounts as follows: they have no prior interaction with the searchers’ contracts and almost all their transactions are related to the attack, often involving a malicious contract created specifically for the exploit.
We applied the above rules to all transactions from 2021 to 2024 to identify real-world attacks on MEV bot contracts. After manually verifying the detected cases, we confirmed a total of 40 attacks targeting MEV bot contracts, involving 21 malicious contracts and 12 victim MEV bot contracts.
The earliest observed attack occurred in July 2021, resulting in a loss of 250 ETH (about \$475K), while the most recent attack happened in October 2024, resulting in a loss of 38 ETH (about \$72K). The total losses from these incidents amount to about \$900K. The losses are calculated using the token price at the time of each attack. We notice that only three of these attacks have been previously reported. It highlights that vulnerabilities in asset management smart contracts remain an ongoing security concern, and often go undetected by existing tools.
Meanwhile, we apply skanf to the victim contracts and find that it successfully identifies all of them as vulnerable. If the searchers had used skanf to secure their contracts at the time, these attacks might have been prevented.
6.2. New Attack Pattern: MEV Phishing Attack
We next delve into the attacks to understand how they occurred. Similar to the “Destroyer Inu” Attack (Sec. 3), in all these incidents, the adversary lures the victim into interacting with their malicious contract by creating a special MEV opportunity. Once searchers try to capture this seemingly profitable MEV through their MEV bot, they fall into the phishing trap, as the malicious contract is now embedded in their MEV supply chain.
We summarize these attacks as MEV phishing attacks and further categorize them into two types based on the attack vectors: token-based and refund-based MEV phishing attacks, comprising 39 and 1 incidents, respectively. A typical example of the former is the “Destroyer Inu” Attack discussed in Sec. 3. Below, we analyze the latter, which introduces a novel approach to MEV phishing.
[…]
Figure 7. The workflow of refund-based MEV phishing attack.
[…]
The attacker first deploys a malicious contract to serve as the refund address (①) and then creates an MEV opportunity, involving a swap, which is submitted to the BackRunMe service (bloXroute Labs, 2025) (②). The tokens used in the swap follow the ERC-20 standard, avoiding detection mechanisms targeting non-compliant tokens.
A searcher receives the opportunity information via BackRunMe (ⓐ) and invokes its MEV bot to extract it (ⓑ). As part of the BackRunMe protocol, the searcher must send a refund to the attacker-specified address (ⓒ). Upon receiving the refund, the malicious contract is triggered and calls the MEV bot contract using crafted calldata (③). Since the transaction originates from the searcher, it bypasses access control and transfers tokens to the attacker.
6.3. Incorrect Access Control vs. No Control
The phishing attacks we analyze bypass incorrect access control mechanisms relying on tx.origin. However, our tool also identifies many contracts that completely lack access control to prevent malicious adversaries.
[…]
We have communicated this issue with the 1inch team. They indicated that the impact is limited, as the contract only contains residual tokens from executed transactions, without affecting any user-owned assets. Interestingly, we observe repeated attempts to extract these residual tokens. Specifically, many attackers exploit the contract whenever its residual token balance exceeds the transaction fee, leaving the contract with little to no available assets. An example of this phenomenon can be found in the transaction with hash 0$\dots$7776660xbfe84e7344f13ac8b9e8c1877bf753862487357cefab64bd555a98edab406777, which earned less than \$$\$1$. In this case, the cost of a rigorous access control mechanism seems less justified, considering that the exploit only extracts residual tokens. It may be more reasonable to save gas for legitimate users of the contract. We discuss the tradeoff between cost and security in Sec. 8.1.
Answer to RQ-3 We identify 40 MEV phishing attacks in the wild, causing about \$900K in losses to searchers. These include 39 token-based attacks and 1 refund-based attack. Only three of them have been covered in prior analyses and reports. skanf detects asset management vulnerabilities in all victim contracts.
7. Related Works
Static analysis.   Static analysis is a method of examining source code or bytecode without executing it to detect vulnerabilities in software. Static analysis tools for smart contracts, such as Vandal (Brent et al., 2018), Securify (Tsankov et al., 2018), MadMax (Grech et al., 2018), Slither (Feist et al., 2019), and Ethainter (Brent et al., 2020), typically aim for completeness but often produce false positives, requiring manual effort to verify warnings. Moreover, a well-known challenge for static analysis is control flow obfuscation (Moser et al., 2007), which may lead tools to struggle to construct accurate control flow graphs, thus reducing their effectiveness. Specifically, potentially vulnerable logic may be treated as unreachable, leading to false negatives. The deobfuscation technique in skanf complements these tools by improving accuracy and efficiency.
Symbolic execution.   Symbolic execution explores a program’s execution paths by treating inputs as symbolic variables rather than concrete values. Symbolic execution tools for smart contracts, such as Oyente (Luu et al., 2016), Mythril (ConsenSys, 2025), teEther (Krupp and Rossow, 2018), Manticore (Mossberg et al., 2019), ETHBMC (Frank et al., 2020), and greed (Gritti et al., 2023; Ruaro et al., 2024), systematically explore possible execution paths to identify issues like reentrancy and integer overflows. However, like static analysis, symbolic execution struggles with symbolic jump destinations caused by control flow obfuscation. skanf extends greed with a de-obfuscation module, allowing it to handle control flow obfuscation and detect vulnerabilities in obfuscated contracts. Although JACKAL (Gritti et al., 2023) and CRUSH (Ruaro et al., 2024) are also built on greed and adopt a similar approach to the symbolic execution used in skanf’s fallback mode, they target different types of vulnerabilities–confused deputy and storage collision, respectively. In contrast, skanf targets asset management vulnerabilities in closed-source contracts and improves performance by guiding concolic execution with historical transactions.
Fuzzing.   Fuzzing is a dynamic testing technique that mutates inputs to uncover software vulnerabilities. For smart contracts, fuzzing applies random and unexpected inputs, such as crafted calldata and reentrant calls, to test contract behavior. Tools developed for EVM smart contracts include ContractFuzzer (Jiang et al., 2018), Echidna (Grieco et al., 2020), Smartian (Choi et al., 2021), Confuzzius (Torres et al., 2021b), ItyFuzz (Shou et al., 2023), and MAU (Chen et al., 2024). However, while state-of-the-art fuzzing tools have improved the mutation efficiency, they remain insufficient for our problem. The vulnerability logic we focus on is often triggered only by specifically crafted calldata, which differs across contracts. As a result, fuzzers may treat these inputs as arbitrary byte mutations, making the exploration process highly inefficient.
Attacks against searchers.   When searchers send transactions to the public mempool, their transactions and execution logic become visible to adversaries, enabling real-time imitation and front-running. Qin et al. (Qin et al., 2023) generalize this attack through Ape, a framework that uses dynamic program analysis to automatically generate adversarial contracts. Interestingly, the same technique can defend against malicious searchers. Zhang et al. propose STING, a defense mechanism that identifies attacking transactions and instantly synthesizes counterattack smart contracts (Zhang et al., 2023).
However, as private mempools become more common (Yaish et al., 2023; Yang et al., 2024a), previous solutions lose effectiveness because searchers’ pending transactions are invisible when sent to a private mempool. Various works attempt to circumvent these limitations. The GhostTX attack by Yaish et al.(Yaish et al., 2024) tricks searchers into bundling adversarial transactions that appear profitable but are invalid, lowering their standing in reputation-based private mempools. Shou et al.(Shou et al., 2024) introduce BACKRUNNER, which exploits (1) the delay between exploit deployment and execution, and (2) incomplete asset drainage, enabling backrunning to recover funds. These works target searchers by analyzing their logic and competing for the same MEV. In contrast, we exploit vulnerabilities within searchers’ contracts directly.
MEV attacks and defenses.   Although MEV searchers are considered victims in this paper, prior work often considers them attackers. They perform various MEV activities to extract profit (Daian et al., 2020; Qin et al., 2022; Li et al., 2023; Ferreira Torres et al., 2024), including front-running (Torres et al., 2021a), sandwiching (Zhou et al., 2021c), liquidations (Qin et al., 2021), and arbitrages (Zhou et al., 2021b; McLaughlin et al., 2023; Heimbach et al., 2024; Öz et al., 2025). Some of these activities, notably front-running and sandwich attacks, constitute direct attacks on users by manipulating their transaction order to extract value. To mitigate the negative effects of MEV on users, both academia and industry have proposed various countermeasures, including private transaction channels (Yang et al., 2024b, a), time-based order fairness protocols (Kelkar et al., 2020; Zhang et al., 2020; Kelkar et al., 2023), and front-running-resistant AMM designs (Zhou et al., 2021a; Wadhwa et al., 2024; Zhang et al., 2024).
8. Discussion
8.1. Tradeoff Between Cost and Security
Unlike traditional programs, computation on EVMs is expensive: users must pay high fees for complex transactions. This implies that, if rigorously enforced, access control can be costly. To illustrate this, we show in Fig. 10 (in Appx. C) how access control for a Uniswap V3 pool requires computing the expected address of the pool according to the CREATE2 specification (Buterin, 2018), consuming 1,534 gas per interaction. Assume that each transaction sent to the MEV bot involves an interaction with a Uniswap V3 pool, requiring one verification. If a searcher sends 10K transactions to the MEV bot contract per month, the total gas usage amounts to 15M. In contrast, if the contract employs an incorrect but cheap verification by checking tx.origin, the total gas usage is reduced to 0.2M.
This raises an interesting question: what is the best tradeoff between cost and security? All MEV bots are driven by economic profits, which means they are extremely sensitive to cost, while a rigorous approach would definitely increase their transaction costs. Therefore, an interesting research direction would be how to provide as many security guarantees as possible while keeping the transaction cost low.
8.2. Implications of EVM Object Format
The EVM Object Format (EOF) (Beregszaszi et al., 2021) is a proposed redesign of smart contract bytecode that introduces structured sections, enforces static control flow, and removes dynamic jumps. These changes aim to improve analyzability and enable future EVM upgrades.
In EOF, control-flow obfuscation becomes significantly more difficult, as indirect jumps and dynamic dispatch are no longer allowed (pcaversaccio et al., 2025). This makes static analysis easier and reduces the viability of the obfuscation techniques targeted in this work.
Nevertheless, the concolic execution approach we present remains effective. Even in the absence of control-flow obfuscation, detecting asset management vulnerabilities requires reasoning about attacker-controlled inputs and complex data flows — tasks that benefit from the symbolic techniques used in skanf.
Moreover, legacy (non-EOF) smart contracts will continue to exist and interact with new contracts, preserving the relevance of our techniques. Until EOF adoption is complete — and perhaps even beyond — tools like skanf will remain essential for analyzing the security of deployed bytecode across formats.
9. Conclusion
We have presented skanf, an EVM bytecode analysis tool optimized for closed-source smart contracts such as MEV bots. skanf can effectively deobfuscate the control flow and identify asset management vulnerabilities, an ability that existing tools do not offer. skanf also employs concolic execution to improve efficacy. We evaluated skanf against real-world MEV bots, using historical transactions to facilitate concolic execution. Among 6,554 MEV bots we studied, skanf detects vulnerabilities in 1,028 of them. Further, skanf automatically generates exploits against 373 of them, with a potential loss exceeding \$9.0M. Furthermore, we discovered 40 attacks in the wild that exploited asset management vulnerabilities against 12 MEV bots, resulting in a total loss of \$900K.
Appendix A Destroyer Inu Attack Analysis
Figure 8. Two scenarios where an MEV bot contract receives a callback: a phishing contract (①) and a DEX (②).
[…]
A direct cause of the Destroyer Inu attack is that the victim contract relies on a vulnerable comparison to block unintended calls from adversaries. A strawman mitigation is to use a hard-coded comparison against msg.sender instead of tx.origin. However, this solution does not meet real-world requirements when verifying whether the caller is the searcher (the creator and owner of this smart contract), because, in practice, the contract may be called by other different contracts. For example, Uniswap V3 pools require any contract that calls their swap function to implement a function called uniswapV3SwapCallback (Uniswap, 2025). This function is invoked by a Uniswap V3 pool at the end of executing swap to ensure that the interacting contract pays the pool the tokens owed for the swap (as illustrated in Fig. 8, scenario ②).
Other DeFi protocols, such as Uniswap V2, Sushiswap, and AAVE, also adopt similar callback designs. In these cases, a hard-coded comparison against msg.sender is not effective because the caller may be any contract deployed by these protocols. Compared to this, verifying tx.origin is not restricted by a specific caller and can prevent direct calls to the contract, but it introduces security risks. Therefore, more precise protection is needed to prevent potential adversaries, which can be a high requirement for contracts, as real-world attack incidents have shown.
Appendix B Historical Versions of MEV Bots
Although smart contracts on Ethereum are generally immutable once deployed, their bytecode at a given address may change over time through the use of SELFDESTRUCT and CREATE2 opcodes (Wood et al., 2014). Specifically, SELFDESTRUCT removes the contract code from the state, and CREATE2 allows redeployment of a new contract at the same address using a fixed deployer address, salt, and modified initialization code. As a result, a smart contract may have different bytecode at different points in time (Ma et al., 2023).
We define a version of a smart contract as the specific bytecode deployed at a given address during a particular period. If a contract is destroyed and re-deployed with different bytecode at the same address, each instance is considered a distinct version.
Version statistics of MEV bots.   To identify all versions of an MEV bot contract, we analyze the historical transactions associated with its address. Specifically, we parse transaction traces to check whether the address was the target of a CREATE or CREATE2 deployment. For each instance of bytecode deployment at the address, we record the corresponding version. This allows us to reconstruct the full sequence of code changes for MEV bot contracts that may have been redeployed multiple times.
We apply this method to all MEV bots in our dataset. As shown in Fig. 9, about 96% of them are deployed only once, indicating no code change over time. Meanwhile, about 0.9% of MEV bots have more than five versions, suggesting that their code has been repeatedly updated.
Figure 9. The distribution of the number of versions per MEV bot in our dataset.
[…]
Vulnerability detection in historical versions.   For MEV bots with multiple historical versions, we also apply skanf to each historical version to identify potential asset management vulnerabilities and attempt exploit generation. This analysis allows us to detect vulnerabilities that may exist only in earlier versions of the MEV bot contract but not in the current one. In total, we identify 114 vulnerable historical versions across 44 MEV bot contracts and successfully generate 41 exploits targeting these versions.
Loss estimation.   To approximate the maximum possible loss, we analyze the highest ERC-20 token balance held by each exploitable version of an MEV bot. Specifically, for each exploitable version, we identify the block range during which that version was active and observe the maximum ERC-20 token balance within that range. If the version is the latest one, we consider the range from its deployment up to block 22,000,000 (corresponding to March 8, 2025).
Consistent with the estimation for current loss, if a contract’s vulnerability is limited to a specific ERC-20 token, we consider only the potential loss associated with that token. If an MEV bot has multiple exploitable versions, we report the loss corresponding to the version with the highest observed ERC-20 balance. To simplify the estimation, we assume that all tokens were sold on March 8, 2025. This allows us to use the token prices on Binance (Binance, 2025) on that day to estimate the potential loss for seven major ERC-20 tokens: WETH, WBTC, USDC, USDT, DAI, UNI and LINK.
Appendix C Solidity Example of Checking Caller
Fig. 10 shows a Solidity example of how to rigorously check whether the caller is a Uniswap V3 pool. The verifyPoolAccess function checks whether the caller is a valid Uniswap V3 pool by reconstructing the expected pool address using the CREATE2 scheme. It first computes a salt by hashing the tuple (token0, token1, fee) with keccak256, consistent with how Uniswap V3 encodes pool parameters. Then, it derives the expected pool address using the standard CREATE2 formula, which combines the factory address, the salt, and the pool’s initialization code hash. If the derived address matches msg.sender, the function returns true, confirming that the caller is a legitimate pool contract created by the Uniswap V3 factory with the given parameters.
1contract UniswapV3PoolAccessControl {
2 address FACTORY = 0x1F98431c8aD98523631AE4a59f267346ea31F984;
3 bytes32 POOL_INIT_CODE_HASH =
4 0xe34f36d28c5efcd7c58e2e84af79e2a dffbe52f705d05dca7e6a181f8a19baf1;
5
6 function verifyPoolAccess(
7 address token0,
8 address token1,
9 uint24 fee
10 ) public view returns (bool) {
11 bytes32 salt = keccak256(abi.encode(token0,token1,fee));
12
13 address computedAddress = address(uint160(uint256(
14 keccak256(abi.encodePacked(
15 bytes1(0xff),
16 FACTORY,
17 salt,
18 POOL_INIT_CODE_HASH
19 )))));
20
21 return computedAddress == msg.sender;
22 }
23}
Figure 10. A Solidity implementation to check if the caller is a Uniswap V3 pool. It verifies whether a given pool address is valid by deriving its expected address using CREATE2.
[…]
A Solidity implementation to check if the caller is a Uniswap V3 pool. It verifies whether a given pool address is valid by deriving its expected address using CREATE2.